OpenStudy (anonymous):

do these lines intersect? [x,y,z]=[2,-3,2]+s[3,4,-10] [x,y,z]=[3,4,-2]+s[-4,5,3]

OpenStudy (anonymous):

Do you want me to show you what I did?

OpenStudy (anonymous):

because I'm not getting the answer my book has

OpenStudy (anonymous):

alright. lets see what you did

OpenStudy (anonymous):

So I made the parametric equations for both lines line 1 : x=2+3s y=-3+4s z=2-10s line 2: x=3-4t y=4+5t z=-2+3t

OpenStudy (anonymous):

Then I set the equations equal to each other. The x of line 1 with the x of like two and so on. I got these 3 equations 3s+4t=1 4s-5t=7 -10s-3t=-4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then I used elimination on 3s+4t=1 and 4s-5t=7. I multiplied 3s+4t=1 by 4 and 4s-5t=7 by 3. I got t=-17/31 after solving

OpenStudy (anonymous):

I then subbed t=-17/31 into 4s-5t=7 to find s. I got s=33/31

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

After getting t and s. I subbed them into -10s-3t=-4 to see if the right side is equal to the left side. I got -9=-4

OpenStudy (anonymous):

but my book has an answer but since I'm given x, y and z. the the point of intersection should also have an x y and z in it. but it gives (33/31, -17/31). But those are my s and t values. so I'm a little confused

OpenStudy (anonymous):

because the question asks if the intersect, and what the points are if they do

OpenStudy (anonymous):

do you think they intersect?

OpenStudy (anonymous):

no, because the left side doesn't equal the right side. Also they are not parallel. So I think they are skew lines. But i think i'm wrong

OpenStudy (anonymous):

alright. i would have done it a little different to you but here is what i did.

OpenStudy (anonymous):

don't both about your s and t values as i chose s and t values for the opposite lines

OpenStudy (anonymous):

so yea i got [3t-4s,4t-5s,-10t-3s]=[1,7,-4]

OpenStudy (anonymous):

ok. now here is the thing

OpenStudy (anonymous):

i'll say it later, let me get it into parametric form.

OpenStudy (anonymous):

so yea your right with parametric form

OpenStudy (anonymous):

3t+4s=1 4t-5s=7 -10t-3s=-4

OpenStudy (anonymous):

for simplicity sake let X represent 3t+4s=1 Y represent 4t-5s=7 and Z yea u get the drift.

OpenStudy (anonymous):

Now for these two lines to intersect, then when solving simultaneously with X and Y these solution should also be the same for when you solve for X and Z or Y and Z You only need to do either 2 solvings. like say X and Y and X and Z or X and Y and Y and Z

OpenStudy (anonymous):

So you have solutions for X and Y which were t=33/31 and s=-17/31

OpenStudy (anonymous):

now for X and Z we get solutions t=13/31 and s=-2/31

OpenStudy (anonymous):

since we get totally different solutions for t and s in both scenario's then these lines do not intersect.

OpenStudy (anonymous):

to check this we can plug these numbers into [3t-4s,4t-5s,-10t-3s]=[1,7,-4] and you will see that you won't get [1,7,-4]

OpenStudy (anonymous):

got any queries?

OpenStudy (anonymous):

so why does my book give me an answer that is the s and t values. Because before in other question it would just say no and that was all

OpenStudy (anonymous):

what do you mean by the s and t values?

OpenStudy (anonymous):

the answer my book gives is (33/31, -17/31)

OpenStudy (anonymous):

the s and t values i get

OpenStudy (anonymous):

thats weird. those values only seem to work of X and Y

OpenStudy (anonymous):

either we are wrong or the book is wrong.

OpenStudy (zarkon):

1) double check that you have the question written correctly...2) if you do then the book is wrong

OpenStudy (anonymous):

my book always numbers the equations and they always use the first two, x and y then sub into z to see if the sides are equal

OpenStudy (anonymous):

i was taught differently. but same concept still applies.

OpenStudy (anonymous):

yea so see if you wrote the question down correctly.

OpenStudy (anonymous):

yea, i wrote is right

OpenStudy (anonymous):

kk. book is silly,

OpenStudy (anonymous):

I want to ask my teacher, but i'm scared to say that the book is wrong and i'm right, since i'm just learning this and i'm not the best person in math

OpenStudy (anonymous):

doesn't hurt to go through your method with your teacher.

OpenStudy (anonymous):

ok. could to stay around. I might have another question

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I have another question .

OpenStudy (anonymous):

My book asks if line r=r(with arrow on top) [-5,1,-2]+k[1,6,5] intersects with the equation [x,y,z]=[2,3,-1]+s[1,3,4]+[-5,4,7]. I know that instead of [x,y,z] i can put r (with arrow on top) but what is with the notation of the first equation, does it mean anything else

OpenStudy (anonymous):

can you draw it for me- just the r arrow thing

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry for my bad drawing

OpenStudy (anonymous):

thats what it has

OpenStudy (anonymous):

thats just a notation for a vector line

OpenStudy (anonymous):

so don't get overwhelmed by it. doesn't contribute to your calculations

OpenStudy (anonymous):

but why does it have r=r with arrow

OpenStudy (anonymous):

i just put it r with arrow = blah blah

OpenStudy (anonymous):

because r itself could be a point but since its r dash then this represents vector notation.

OpenStudy (anonymous):

ok thank you so much for your help

OpenStudy (zarkon):

You can write vectors on the site \[\vec{r}\] \vec{r}